A football is launched at an angle such that the vertical component of its velocity and the horizontal component of its velocity are both equal to 40 m/s.
How long does it take for the football to reach its highest point?
What horizontal distance does the football travel in this time?
1
Expert's answer
2019-10-21T09:05:12-0400
"t=\frac{v}{g}=\frac{40}{9.8}=4.1\ s"

"x=vt=\frac{v^2}{g}=\frac{40^2}{9.8}=160\ m"